Bob's father is 3 times old as Bob.4 years ago,he was 4 times older.How old is Bob?
Answer by uday1435(57)

Let Bob's age be ‘b’ and is father's age ‘f’
3b = f
Four years ago Bob was (b – 4) years and his father was (f – 4) years
4(b – 4) = 4(f – 4) but f = 3b
So 4( b – 4) = (3b – 4)
4b – 16 = 3b – 4
So b = 16 – 4 = 12 and f = 3b = 3(12) = 36
Answer Bob is 12 years old and his father is 36 years old





If you still require further clarifications please write to udayakumar.t.r@gmail.com